jimmy has a large house, he needs to fence his yard so his dog can run and play without running away. The yard is 20.5 feet by 16.2 feet. the fencing costs $9.50 for 3 yards. how much will it cost him to fence the whole yard? (round your answer to the nearest cent)
Answer by checkley77(12844)

20.5*16.2=332.1 FT. OF FENCING IS NEEDED.
(332.1/9)*9.50=36.9*9.50=$350.55 IS THE COST OF THE FENCE.